# Karma: An Experimental Study

### v1

@@ 2.3. Experimental Implementation
Table 3 in Appendix A lists all parameter values used in the experiments.
3. Results
@@
3.1. Welfare Measures
Our central welfare measure is the efficiency gain which we define next. For a particular participant $i$, let $\left(u_{i}(t)\right)_{t\in\{1,\dots,T\}}$ be the vector of realized urgency in the experiment, $S_{i}=s_{i}(T)$ the total score at the end of the experiment, and $S^{\textup{rand}}_{i}=\frac{1}{2}\sum_{t=1}^{T}u_{i}(t)$ the ex-ante expected total score, given the urgency realization, under random allocation. Then the efficiency gain is defined as
(1) $E_{i}=\frac{S_{i}-S^{\textup{rand}}_{i}}{S^{\textup{rand}}_{i}},$
and gives the relative improvement of participant $i$ with respect to the ex-ante expected random allocation. On this basis, we will assess efficiency based on the median efficiency gain among participants, and fairness based on the distribution of efficiency gains.
3.2. High-level Synopsis of Results
Overall, we find that there are pronounced and statistically significant efficiency gains in all treatments, higher under the high stake process than under the low stake process, but not significantly different from one another in terms of overall efficiency gains depending on whether the bidding scheme is binary or full range. In all treatments, more than $90\%$ of the population is better off with karma than under random allocation. The most favorable combination is high stake with binary bidding, both in terms of median efficiency gains and in terms of the distribution of gains as most individuals are on the benefiting side.
Remark: With the best intentions, we had fully pre-registered design and analysis of our experiments. Some but not all of the analyses presented here were indeed pre-registered, and there are also further analyses from the pre-registration that are not presented in this paper. We stuck with the pre-registration as much as we could, but the unfortunate need to depart from the pre-registered plan of analysis is explained in Appendix B.
3.3. Efficiency Results
Efficiency went up in all treatments. Figure 2 shows the median efficiency gain in each of the four treatments (denoted ‘total’), cf. Table 1, as well the median efficiency gain in the bottom and top halves of the population in terms of individual outcomes. For the purpose of visualizing the data spread, estimates of the $95\%$ confidence intervals (CI) are also shown in the figure based on 1000 bootstraps. [^1] Note that data samples are dependent for participants in the same experiment, therefore the bootstrapped CIs should be seen as heuristic estimates and for visualization purposes. For the random allocations, the CI was estimated by running 1000 independent simulations.
@@ 3.3. Efficiency Results
Figure 2. Median efficiency gains for the four treatment combinations, cf. Table 1, with estimate of $95\%$ confidence interval. Also shown are the medians in the lower and upper halves of the population.
Table 2. Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains under karma versus random for all treatments. Karma treatments have $n_{\textup{karma}}=100$, and random is based on $n_{\textup{random}}=100\,000$ samples.
@@
3.4. Fairness Results
Figure 2 shows that in addition to achieving overall efficiency gains, the karma allocation is more efficient than random allocation for both the lower/less fortunate and the higher/more fortunate halves of the population. To provide finer grained insight on the fairness of karma, Figure 3 further shows the mean efficiency gain per population decile for the four treatments. It is important to note that while the efficiency gain (1) is defined with respect to the ex-ante expected score under random allocation, not the whole population realizes this score ex-post. To account for the fact that there will be more or less fortunate individuals under random allocation, Figure 3 also shows the ex-post mean efficiency gains per population decile for random allocation, with the $95\%$ confidence interval estimated from 1000 independent simulations per treatment. The key feature to observe is that in all treatments, $90\%$ of the population achieve higher efficiency in the karma scheme than the random scheme. Only the lowest decile is worse-off in karma than random. This decile is dominated by non-adopters whose bids defaulted to zero due to inactivity (number of non-adopters: low stake-binary: 9/100; low stake-full range: 6/100; high stake-binary: 7/100; high stake-full range: 5/100). Non-adopters achieve particularly low scores as with a consistent bid of zero it is very unlikely to get granted priority.
Figure 3. Mean efficiency gain per decile for the four treatment combinations.
Moreover, it is evident that the treatment combination of high stake-binary is particularly favorable. Not only does it lead to the highest total median efficiency gain, cf. Figure 2, but it also achieves the largest gap to random allocation across deciles, cf. right panel of Figure 3. This suggests that all (active) subgroups of the population achieve strong benefits in this treatment.
@@ 3.4. Fairness Results
4. Discussion
@@
[…]
Another way to interpret the observed differences between online subjects and Nash or expert play may be related to impatience or perception of future rounds as less important. To investigate this, we compared the aggregate efficiency gains realized in our online experiments to those that would be achieved in theory in a short-sighted equilibrium, as predicted by (Elokda et al., 2023). It was found that the realized gains of online subjects correspond to if players discount future rounds by a factor of approximately 0.3–0.7, while those of experts correspond to a future discount factor of more than 0.98.
Another important consideration regarding implementation of karma with human participants is whether a simpler scheme (binary) or a richer scheme (full range) is favourable. That fact that we found no significant differences in terms of realized efficiency gains between the two provides preliminary evidence that the simpler binary scheme–at least for applications similar to the ones we studied–can be employed without loss in efficiency. Such applications feature, as we have investigated in our experiments, urgency processes with binary levels, for which a binary scheme is arguably also particularly natural. Theoretically, Nash equilibrium under binary bidding will lose in efficiency with more than two urgency levels, and it remains an open question to test what the trade-off is between behavioral simplicity of a binary (or otherwise limited) bidding scheme and the theoretical benefits that come with richer schemes. Some reasons to believe simplicity is beneficial are based on theories of decision fatigue (Baumeister, 2003; Pignatiello et al., 2020) and simplicity in mechanism design (Pycia and Troyan, 2023). A richer analysis of this complexity-performance trade-off provides fruitful avenues for future investigations, both in theory and in the behavioral lab.
@@ Appendix A Supplementary Implementation Details
Figure 4. Decision page examples.
Figure 5. Results page example.
Table 3. List of notation and parameter values
@@ Appendix B Remarks regarding departure from pre-registered analysis plan
Originally, we pre-registered the design of the experiment along with a complete plan for analysis at the Open Science Framework under [link withheld to preserve anonymity of submission]. The analysis presented in the paper above constitutes a deviation from this pre-registered analysis plan. Unfortunately, we had to re-design our analysis and statistical tests due to a technical bug in the logging of a participant’s bid who made a non-zero bid but was outbid (the ‘loser’ that round) by the other participant (the ‘winner’ that round), and therefore did not spend any karma. In scenarios like this, the winner and the karma accounting were done correctly, but a bug in the data-logging implementation resulted in a loss of data concerning the bid levels of the losers, thus making them indistinguishable also from players who bid zero. This loss of data logging made it impossible to perform the pre-registered analysis, and we opted for the alternative analysis presented above. Here, for completeness, we state the hypotheses that were originally pre-registered:
  1. (1)
@@
As regards the efficiency-related hypotheses (1), (3), (6) and (8), the results of our MWW tests strongly support hypothesis (1), weakly support hypothesis (8), and reject hypotheses (3) and (6). Note, however, that our original plan was to measure realized efficiencies using ex-post means, which was found to be overly sensitive to stochastic effects. Therefore, the analysis presented in Section 3 is based on ex-ante expected medians instead. As regards the fairness-related hypotheses (2), (4) and (7), our fairness analysis supports hypotheses (2) and (7), in particular for the high stake treatments, but not hypothesis (4). These conclusions are drawn based on comparisons of the efficiency gains across deciles, and not as measured by the negated standard deviations of final scores, as we had originally planned, which turned out to be a measure that does not produce interpretable results. Unfortunately, due to the bug, hypothesis (5) could not be tested.
@@ Appendix C Supplementary Results
Table 4. Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains in the first versus second halves of the experiments for all treatments. All treatments have $n_{\textup{karma}}=100$ samples.
@@
Table 5. Inter-treatment efficiency gain MWW tests reporting test statistics $U$ (with associated $p$-values in brackets). All treatments have $n=100$ samples.
@@
Binary
Full range
Combined
Binary
Full range
Combined Binary Full range
@@
Figure 6 compares the median efficiency gains realized in our online MTurk experiments with a) the gains that would have been achieved if the participants followed the Nash equilibrium (as determined by running 1000 simulations), and b) for the low stake-full range treatment, the gains achieved by ‘expert’ subjects, those are, students of an applied game theory class [^4] The median efficiency gain of Expert is higher than Nash due to a large number of inactive participants (9/28). Inactive participants were added due to a technical issue in forming an experimental group that required setting the size of the group greater than the number of students.. This suggests that stronger benefits can be achieved with more training and commitment.
Figure 6. Comparison of the median efficiency gains in the online MTurk experiments versus simulations of the Nash equilibrium. For the low stake-binary treatment, the median efficiency gain from an ‘expert’ game theory class experiment is also shown.

### v2

@@ 3.3 Experimental Implementation
Table 3 in A lists all parameter values used in the experiments.
4 Results
@@
4.1 Welfare Measures
Our central welfare measure is the efficiency gain which we define next. For a particular participant $i$, let $\left(u_{i}(t)\right)_{t\in\{1,\dots,T\}}$ be the vector of realized urgency in the experiment, $S_{i}=s_{i}(T)$ the total score at the end of the experiment, and $S^{\textup{rand}}_{i}=\frac{1}{2}\sum_{t=1}^{T}u_{i}(t)$ the expected total score, given the urgency realization, under random allocation. Then the efficiency gain of participant $i$ is defined as
$E_{i}=\frac{S_{i}-S^{\textup{rand}}_{i}}{S^{\textup{rand}}_{i}}.$ (1)
This definition expresses the relative improvement with respect to the expected random score given the urgency realization, in order to control for randomness in the urgency realization. On this basis, we will assess overall efficiency based on the median efficiency gain among participants, and fairness based on the distribution of efficiency gains.
@@ 4.1 Welfare Measures
4.2 High-level Synopsis of Results
Overall, we find that there are pronounced and statistically significant efficiency gains in all treatments, higher under the high stake process than under the low stake process, but not significantly different from one another in terms of overall efficiency gains depending on whether the bidding scheme is binary or full range. In all treatments, more than $90\%$ of the population is better off with karma than under random allocation. The most favorable combination is high stake with binary bidding, both in terms of median efficiency gains and in terms of the distribution of gains as most individuals achieve pronounced benefits.
Remark: With the best intentions, we had fully pre-registered design and analysis of our experiments. Some but not all of the analyses presented here were indeed pre-registered, and there are also further analyses from the pre-registration that are not presented in this paper. We stuck with the pre-registration as much as we could, but the unfortunate need to depart from the pre-registered plan of analysis is explained in B.
4.3 Efficiency Results
Efficiency went up in all treatments. Figure 2 shows the median efficiency gain in each of the four treatments (denoted “Total”), cf. Table 1, as well the median efficiency gain in the bottom and top halves of the population in terms of individual outcomes, contrasted to the median gains that would be realized if all participants followed theoretically optimal Nash behavior. For the purpose of visualizing the data spread, estimates of the $95\%$ confidence intervals (CI) are also shown in the figure based on 1000 bootstraps [^3] Note that data samples are dependent for participants in the same experiment, therefore the bootstrapped CIs should be seen as heuristic estimates for visualization purposes. Similar estimates were attained for larger numbers of bootstraps.. For the random and Nash allocations, the CI was estimated by running 1000 independent simulations.
The findings of Figure 2 summarize as follows. Although the experimental efficiency gains fall short of the theoretical Nash gains, there are nonetheless positive total efficiency gains in all treatments (median ranging from $7.38\%$ to $15.34\%$). The efficiency gains are especially pronounced in the upper half of the population (median ranging from $22.51\%$ to $29.85\%$); whereas the lower half of the population does not perform much worse than ex-ante expected under random allocation (median ranging from $-6.5\%$ to $-0.65\%$), and in fact performs better than the lower half of the population under random allocation ex-post. This suggests that overall, karma provides an opportunity for participants to achieve pronounced benefits, without strongly harming those that are less strategic. Moreover, the efficiency gains are statistically significant in all treatments, as elaborated in Table 2 reporting MWW test statistics $U$ (with associated $p$-values).
Figure 2: Median efficiency gains for the four treatment combinations, cf. Table 1, with estimate of $95\%$ confidence interval. Also shown are the medians in the lower and upper halves of the population.
Table 2: Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains under karma versus random for all treatments. Karma treatments have $n_{\textup{karma}}=100$, and random is based on $n_{\textup{random}}=100\,000$ samples.
@@ 4.3 Efficiency Results
4.4 Fairness Results
Figure 2 shows that in addition to achieving overall efficiency gains, the karma allocation is more efficient than random allocation for both the lower/less fortunate and the upper/more fortunate halves of the population. To provide finer grained insight on the fairness of karma, Figure 3 further shows the mean efficiency gain per population decile for the four treatments. It is important to note that while the efficiency gain (1) is defined with respect to the ex-ante expected score under random allocation given the urgency realization, not the whole population realizes this score ex-post. To control for the fact that there will be more or less fortunate individuals due to radnomness under random allocation, Figure 3 also shows the ex-post mean efficiency gains per population decile for random allocation, with the $95\%$ confidence interval estimated from 1000 independent simulations per treatment. The key feature to observe is that in all treatments, $90\%$ of the population achieve higher efficiency in the karma scheme than the random scheme. Only the lowest decile is worse-off in karma than random. This decile is dominated by non-adopters whose bids defaulted to zero due to inactivity (number of non-adopters: low stake-binary: 9/100; low stake-full range: 6/100; high stake-binary: 7/100; high stake-full range: 5/100). Non-adopters achieve particularly low scores as with a consistent bid of zero it is very unlikely to get granted priority.
Figure 3: Mean efficiency gain per decile for the four treatment combinations.
Moreover, it is evident that the treatment combination of high stake-binary is particularly favorable. Not only does it lead to the highest total median efficiency gain, cf. Figure 2, but it also achieves the largest gap to random allocation across deciles, cf. right panel of Figure 3. This suggests that all (active) subgroups of the population achieve pronounced benefits in this treatment.
@@ 4.4 Fairness Results
4.5 Analysis of Bidding Behaviors
[…]
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 4: Median bid per urgency and karma for the four treatment combinations.
The main finding of Figure 4 is that overall, there was a tendency to over-bid in the low urgency state [^4] Notice that due to the aforementioned bug which excluded some of the losing bids, the medians in Figure 4 are potentially biased towards higher winning bids. Nonetheless, there are sufficiently many samples to support the observed tendency to over-bid in low urgency.. The bidding behaviors are not well-explained by a single value of future discounting: the median bids are well-fitted to far-sighted behavior (discount factors 0.80.80.80.80.980.980.980.98) in high urgency, but to short-sighted behavior (discount factor $0.6$) in low urgency. In contrast, the median bids of top-performing participants are well-fitted to far-sighted behavior in both urgency levels, which explains the high efficiency gains achieved by these participants.
To provide further insight on how over/under-bidding during low/high urgency affected performance, Figure 5 shows a scatter of the mean signed difference to Nash-$0.98$ bids per participant and urgency level, versus the attained efficiency gains, for the four treatments. The data points associated to the top-performing decile are highlighted in orange. As expected, under-bidding in high urgency negatively affects performance, cf. third quadrant of the high urgency panels, which includes the majority of participants achieving negative efficiency gains. On the other hand, the effect of over-bidding in either of the urgency levels is inconclusive. Consistently in all treatments, the single-top performing participant bid close to Nash [^5] The mean absolute difference to Nash-$0.98$ bids correlates well with the efficiency gain, as expected, and particularly so in the full-range treatments (Spearman correlations: low stake, binary: $-0.345$, $p=0.0004$; low stake, full-range: $-0.664$, $p=0.0000$; high stake, binary: $-0.119$, $p=0.236$; high stake, full-range: $-0.674$, $p=0.0000$). However, several top performing participants also tended to over-bid in low urgency, and a few of these participants managed to over-bid in both low and high urgency, cf. Figure 5(d).
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 5: Efficiency gain vs. mean difference to Nash for the four treatment combinations.
4.6 Analysis of stationarity
An important feature of the karma mechanism is that it forms a closed economy in which karma is preserved over time. In theory, this feature enables reaching a stationary regime, in which a predictable, time-invariant optimal behavior can be repeated indefinitely. To investigate whether stationarity is achieved in practice, each sub-figure of Figure 6 shows the variations in the distribution of karma (left panel) and bids (right panel) over the course of the main experimental rounds, for each of the four treatments. The variation in the distributions is measured by the Wasserstein-1 distance, also known as the Earth mover’s distance, between the realized karma/bid distributions of successive rounds. The choice of the Wasserstein-1 distance is interpretable: for example, a distance of one corresponds to shifting the distribution by one karma unit on average. For each treatment, the realized distribution variations are plotted for each of the five experimental groups (labelled “MTurk-G1–G5”), and contrasted to the variations that would be attained if all participants followed stationary Nash behavior (labelled “Nash”; the shaded area coincides with the 95% confidence interval estimated from 1000 independent simulations per group).
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 6: Variation in the karma and bid distributions over time for the four treatment combinations.
The main finding of Figure 6 is that in all treatment combinations, after a short initial transient of less than five rounds, the realized variations in the karma and bid distributions are close in magnitude to those attained under stationary Nash play, as visualized by the shaded Nash area subsuming most of the experimental data. Two exceptions are the bid distributions in the full range treatments, cf. right panels of Figures 6(b), 6(d), in which the experimental variations lie at a notably higher magnitude than the stationary Nash variations. Nonetheless, these variations are still relatively small and oscillate around a Wasserstein-1 distance of one. Overall, this finding suggests that despite the presence of noisy bids, the karma auctions quickly become predictable over time, and these auctions are even more predictable in the binary treatments.
5 Discussion
@@
It is noteworthy that our experimental subjects were recruited from a population of totally untrained and inexperienced users from an online convenience sample (on MTurk). Their behavior, even though more efficient than random, however, is not as efficient as is theoretically feasible under Nash equilibrium play. Analysis of the experimental bidding behaviors reveals a consistent tendency to over-bid in low urgency rounds, however, this observation does not explain the gap to Nash efficiency entirely, which is likely attributed to irrational, noisy behavior. A natural follow-up question is, therefore, whether karma could be capable of achieving higher efficiency gains than realized in our online experiments if the human population consisted of participants that were better trained. As a first step in this direction, we conducted an auxiliary experiment under the low stake-full range treatment with a group of ‘expert’ subjects, who were graduate students in an applied game theory class. It turned out that, indeed, the achieved efficiency gains of these experts are close to Nash levels (MTurk: median $12.86\%$, $n=100$; experts: median $36.65\%$, $n=28$, MWW test $U=667.0$, $p=0.0000$). Therefore, we may interpret the efficiency gains that were realized in our online experiment as behaviorally robust lower bounds on the performance of the karma scheme given the relatively low training and commitment of the subject pool considered.
Another important consideration regarding implementation of karma with human participants is whether a simpler scheme (binary) or a richer scheme (full range) is favourable. That fact that binary led to less variations in the distribution of bids and thus more predictable auction outcomes, meanwhile we found no significant differences in terms of realized efficiency gains compared to full range, provides preliminary evidence that the simpler binary scheme is advantageous–at least for applications similar to the ones we studied. Such applications feature, as we have investigated in our experiments, urgency processes with binary levels, for which a binary scheme is arguably also particularly natural. Theoretically, Nash equilibrium under binary bidding will lose in efficiency with more than two urgency levels, and it remains an open question to test what the trade-off is between behavioral simplicity of a binary (or otherwise limited) bidding scheme and the theoretical benefits that come with richer schemes. Some additional reasons to believe simplicity is beneficial are based on theories of decision fatigue (Baumeister, 2003; Pignatiello et al., 2020) and simplicity in mechanism design (Pycia and Troyan, 2023). A richer analysis of this complexity-performance trade-off provides fruitful avenues for future investigations, both in theory and in the behavioral lab.
@@ Appendix A Supplementary Implementation Details
Figure 7: Decision page examples.
Figure 8: Results page example.
Table 3: List of notation and parameter values
@@ Appendix B Remarks regarding departure from pre-registered analysis plan
Originally, we pre-registered the design of the experiment along with a complete plan for analysis at the Open Science Framework (Elokda and Nax, 2023). The analysis presented in the paper above constitutes a deviation from this pre-registered analysis plan. Here, for completeness, we state the hypotheses that were originally pre-registered, and discuss our motivations for deviating from the pre-registered plan.
  1. 1.
@@
As regards the efficiency-related hypotheses (1), (3), (6) and (8), the results of our MWW tests strongly support hypothesis (1), weakly support hypothesis (8), and reject hypotheses (3) and (6). Note, however, that our original plan was to measure realized efficiencies using ex-post means, which was found to be overly sensitive to stochastic effects. Therefore, the analysis presented in Section 4 is based on ex-ante expected medians instead. As regards the fairness-related hypotheses (2), (4) and (7), our fairness analysis supports hypotheses (2) and (7), in particular for the high stake treatments, but not hypothesis (4). These conclusions are drawn based on comparisons of the efficiency gains across deciles, and not as measured by the negated standard deviations of final scores, as we had originally planned, which turned out to be a measure that does not produce interpretable results. Finally, the bidding behaviors portrayed in Figure 4 weakly support hypothesis (5), however, due to the bug in the logging of bids discussed in Section 4.5, we could not rigorously test this hypothesis.
@@ Appendix C Supplementary Results
Table 4: Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains in the first versus second halves of the experiments for all treatments. All treatments have $n_{\textup{karma}}=100$ samples.
@@
Table 5: Inter-treatment efficiency gain MWW tests reporting test statistics $U$ (with associated $p$-values in brackets). All treatments have $n=100$ samples.
@@
  Binary   Full range Combined   Binary   Full range Combined   Binary   Full range